# SIM-Shapley: A Stable and Computationally Efficient Approach to Shapley Value Approximation

### v1

@@ Abstract
Explainable artificial intelligence (XAI) is essential for trustworthy machine learning (ML), particularly in high-stakes domains such as healthcare and finance. Shapley value (SV) methods provide a principled framework for feature attribution in complex models but incur high computational costs, limiting their scalability in high-dimensional settings. We propose Stochastic Iterative Momentum for Shapley Value Approximation (SIM-Shapley), a stable and efficient SV approximation method inspired by stochastic optimization. We analyze variance theoretically, prove linear $Q$-convergence, and demonstrate improved empirical stability and low bias in practice on real-world datasets. In our numerical experiments, SIM-Shapley reduces computation time by up to 85% relative to state-of-the-art baselines while maintaining comparable feature attribution quality. Beyond feature attribution, our stochastic mini-batch iterative framework extends naturally to a broader class of sample average approximation problems, offering a new avenue for improving computational efficiency with stability guarantees. Code is publicly available at https://github.com/nliulab/SIM-Shapley.
1 Introduction
As machine learning (ML) and artificial intelligence (AI) models grow increasingly complex, explainable AI (XAI) has emerged as a critical area for understanding model behavior and building trust in predictions doshivelez2017rigorousscienceinterpretablemachine ; covert2020understandingglobalfeaturecontributions . Quantifying feature importance is therefore essential in scientific and engineering domains where explainability is non-negotiable. One such domain is healthcare, where understanding the influence of individual predictors can inform targeted interventions and resource allocation ning2024variable .
The Shapley Value (SV) P-295 , originating in cooperative game theory, has become a foundation for XAI because it attributes a feature’s contribution to predictive performance while satisfying key fairness axioms lundberg2017unifiedapproachinterpretingmodel ; rodemann2024explainingbayesianoptimizationshapley ; 7546525 . SV-based explanations also support contrastive explanations, identify corrupted or influential features, and often align well with human intuition covert2020understandingglobalfeaturecontributions ; molnar2025 .
However, they incur exponential computational cost. Let $D=\{1,\dots,d\}$ be the index set of features and $S\subseteq D$ be a subset. A cooperative game is a set function $v:2^{d}\to\mathbb{R}$ that assigns a score to every subset $S$. To obtain the SV of feature $i\in D$, one must average its marginal contribution $v(S\cup\{i\})-v(S)$ over all $S$ that exclude $i$. Because there are $2^{d-1}$ such subsets for each feature, the total number of evaluations grow as $\mathcal{O}(d2^{d})$, which is impractical for large $d$.
To overcome this burden, we reformulate the problem as a stochastic optimization problem AguidetoSAA ; hannah2015stochastic . Our method, SIM-Shapley, iteratively refine estimates via mini-batch sampling and momentum updates polyakAverage ; RuppertAverage ; kingma2017adammethodstochasticoptimization ; tieleman2012rmsprop . An $\ell_{2}$ penalty ridge further reduces variance and enhances stability, yielding a flexible iterative framework broadly applicable to sample-average-approximation (SAA) problems. Our main contributions include:
1. A new SV estimator with provable linear $Q$-convergence, providing faster convergence than state-of-the-art methods while retaining consistency.
2. Stability mechanisms: $\ell_{2}$ regulation, detection of negative-sampling events, and correction of initialization bias.
3. A general, scalable framework compatible with diverse cooperative game formulations, naturally integrating techniques such as paired sampling covert2021improvingkernelshappracticalshapley , and extensible to a wider class of SAA objectives.
2 Related Work
2.1 Shapley Value as Weighted Least Squares
@@
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ denote $d$ features for $n$ samples, and let $D=\{1,2,\dots,d\}$ represent their index set. Given a predictive model $f$ for the response $Y$, the performance of any subset $S\subseteq D$ can be expressed through a cooperative game $v(S)$. A restricted model $f_{S}=f(x_{S},X_{S^{c}})$ uses only features in $S$, with the complement $S^{c}=D\setminus S$ drawn from the conditional distribution $p(X_{S^{c}}\mid X_{S}=x_{S})$, yielding local explanations.
@@
Several choices for $v(S)$ exist. SHAP lundberg2017unifiedapproachinterpretingmodel defines $v(S)$ as the expectation of the prediction from $f_{S}$, whereas Shapley Effects sveffects use output variance. In this work, we adopt the prediction-loss game:
@@
following SAGE covert2020understandingglobalfeaturecontributions . Besides local explanations, global explanations arise by treating an input-label pair $U=(X,Y)$ as exogenous covert2021improvingkernelshappracticalshapley and defining a stochastic cooperative game as:
@@
which captures the expected contribution of a feature subset across the entire data distribution covert2020understandingglobalfeaturecontributions . This formulation provides a unified framework that encompasses both local and global Shapley value–based methods.
The Shapley value P-295 (SV) $\phi_{i}$ of feature $i\in D$ is defined as the average marginal contribution of feature $i$ across all subsets $S$ that exclude $i$:
@@
Lundberg et al. lundberg2017unifiedapproachinterpretingmodel recast Shapley estimation as kernel-weighted linear regression, later refined by Covert et al.  covert2021improvingkernelshappracticalshapley . Approximating $v(S)$ with the additive form $u(S)=\beta_{0}+\sum_{i\in S}\beta_{i}$ (where $\beta_{0}=v(\emptyset),\sum_{i\in S}\beta_{i}=v(S)-v(\emptyset)$) leads to the weighted least squares objective:
@@
where $\mu_{\text{Sh}}(S)$ is the Shapley kernel that yields the optimal solution equal to SV Charnes1988 ; lundberg2017unifiedapproachinterpretingmodel .
Let $\bm{\beta}=(\beta_{1},\dots,\beta_{d})$ denote the optimal coefficients (i.e., the minimizer) for Eq. (4); these are the estimated SV. To represent subsets more conveniently, map each subset $S\subseteq D$ to a binary vector $z\in\{0,1\}^{d}$, with $z_{i}=1\ \Leftrightarrow i\in S$. With a slight abuse of notation, write $v(z)=v(S)$ and $\mu_{\text{Sh}}(z)=\mu_{\text{Sh}}(S)$ for the subset $S=\{i:z_{i}=1\}$. We then define a sampling distribution $p(z)\propto\mu_{\text{Sh}}(z)$ over binary vectors $z$ satisfying $0<\mathbf{1}^{\top}z<d$; otherwise, we set $p(z)=0$.
Under this formulation, estimating SV reduces to solving the following constrained optimization problem covert2021improvingkernelshappracticalshapley :
@@
2.2 Monte Carlo Estimation and Practical Challenges
Evaluating Eq. (5) on all $2^{d}$ subsets is computationally prohibitive. Instead, we draw $m$ i.i.d. samples $z_{i}\sim p(z)$ and their values $v(z_{i})$, yielding the Monte-Carlo surrogate:
@@ 2.2 Monte Carlo Estimation and Practical Challenges
Covert et al. covert2021improvingkernelshappracticalshapley derive a closed-form solution via the Karush–Kuhn–Tucker boyd2004convex (KKT) conditions and proved consistency as $m\to\infty$. In practice, however, only a finite number of $z_{i}$ are available, which may lead to a singular system in certain regions of the solution space and cause the optimization to collapse. Moreover, naively averaging batch estimates yields high variance and fluctuation due to insufficient coverage of the subset space, leading to numerical instability and optimization failure.
2.3 From Sample Average Approximation to Stochastic Iteration
The constrained problem in Eq. (5) and its sampled version in Eq. (6) follow the Sample Average Approximation (SAA) paradigm AguidetoSAA ; shapiro1996convergence : we replace the expectation in Eq. (5) by the empirical average in Eq. (6). Under standard convexity and regularity conditions, the SAA solution converges almost surely as $m\to\infty$ fu2006gradient ; AguidetoSAA . However, large $m$ can be impractical due to computational constraints, and not all samples are equally informative for improving the estimator.
To address this, we turn to Stochastic Approximation (SA) methods hannah2015stochastic ; shapiro1996simulation , which update the solution incrementally using only a mini-batch at each step. This avoids repeatedly solving the full SAA problem and can lead to faster convergence in practice. As pointed out by Royset and Szechtman RePEc:inm:oropre:v:61:y:2013:i:3:p:762-776 , SA can outperform SAA in convergence rate, regardless of whether the underlying solver is sublinear, linear, or superlinear.
Moreover, sampling mini-batches of $z_{i}$ from the $2^{d}$ feature space inflates variance slows convergence qian2020impactminibatchsizevariance ; balles2017couplingadaptivebatchsizes . Adding an $\ell_{2}$ penalty stabilizes updates by shrinking large fluctuations in the estimator, reducing variance without breaking convexity fan2021communicationefficientaccuratestatisticalestimation ; shamir2014communicationefficientdistributedoptimization .
In Section 3, we introduce SIM-Shapley which combines stochastic momentum, adaptive averaging, and $\ell_{2}$ regularization to accelerate convergence and enhance numerical stability in SV estimation.
@@ 3.1 Stochastic Iteration with Momentum
Following the stochastic optimization formulation discussed in Section 2.3, we reformulate the KernelSHAP approximation lundberg2017unifiedapproachinterpretingmodel in Eq. (6) as an iterative procedure to update estimates $\bm{\beta}^{(n)}$ using an exponential moving average (EMA) scheme:
@@
where $10<t<$ 1 is a fixed momentum parameter. $\bm{\delta}^{(n)}$ represents the sampling information obtained at the $n$-th iteration, which can be interpreted as a form of momentum. The updated estimator $\bm{\beta}^{(n)}$ is constructed as a weighted combination of the previous estimate $\bm{\beta}^{(n-1)}$ and $\bm{\delta}^{(n)}$. In addition, $\bm{\beta}^{(n)}$ shall satisfy the constraint in Eq. (5) to preserve the additive property of SV, as reflected in the definition of $u(S)$ (Eq. (4)).
@@
For notational simplicity, we do not distinguish between $v(S)$ in Eq. (1) and $\mathbf{V}(S)$ in Eq. (2), as replacing $v(S)$ with $\mathbf{V}(S)$ allows $\bm{\beta}^{(n+1)}$ in SIM-Shapley to be interpreted as an estimator of the global SV explanation. To approximate $\mathbf{V}(S)$, we construct a Global Reference Set by sampling from the original dataset, thereby approximating the distribution $p(U)$ and enabling the evaluation of $\mathbb{E}_{U}[v(S,U)]$ for each cooperative game.
As discussed in Section 2.1, to approximate the conditional distribution $p(X_{S^{c}}\mid X_{S}=x_{S})$, we handle the missing features $X_{S^{c}}$ by marginalizing them using their joint marginal distribution. Specifically, we build a Background Set from the data and use it to impute the removed variables. Since our focus is on estimating the SV rather than the underlying games themselves, this treatment does not affect the convergence or correctness of the iterative procedure.
Building on these ideas, we propose the SIM-Shapley algorithm, together with its local explanation version described in Algorithm 1. The global version is provided in the Appendix B.
Next we apply the Karush–Kuhn–Tucker (KKT) conditions boyd2004convex to Eq. (7), leveraging the convexity of the objective function and the linearity of the constraint. Given a sampled batch $(z_{1},\dots,z_{m})$, the associated Lagrangian function, with Lagrange multiplier $\nu\in\mathbb{R}$, is defined as follows:
@@ 3.2 Variance and Convergence
We further analyze the variance and convergence behavior of the SIM-Shapley estimator $\bm{\beta}^{(n)}$ with respect to the ideal Shapley value vector $\bm{\beta}$. Throughout this section, we treat the KernelSHAP estimator in Eq. (6) as the ground truth, consistent with prior work demonstrating that KernelSHAP and its variants are either unbiased or nearly unbiased in practice covert2021improvingkernelshappracticalshapley ; covert2020understandingglobalfeaturecontributions ; chen2022algorithmsestimateshapleyvalue ; lundberg2017unifiedapproachinterpretingmodel . Since SIM-Shapley is designed to approximate KernelSHAP through a stochastic iterative scheme with regularization, any potential bias is empirically negligible, as confirmed by our experiments.
@@
In the following, we focus on the variance and convergence properties of SIM-Shapley across iterations, as they dominate the practical performance in typical settings where bias is negligible as expected, as later shown in Section 5.
@@
Let $\hat{\bm{\beta}}$ denote the estimator obtained from KernelSHAP in Eq. (6) with $m$ samples of $z$, and let $\bm{\beta}^{(n)}$ be the SIM-Shapley iterates with fixed momentum $t\in(0,1)$. As $m\to\infty$, for fixed momentum $t\in(0,1)$ and any feature index $i$ and any iteration $n\geq 1$,
@@
A detailed proof of Theorem 1 is given in Appendix A.1. In essence, each EMA update in SIM-Shapley reduces the variance of feature $i$ by at least a factor of $(1-t)^{2}$ relative to the vanilla SAA estimator $\hat{\bm{\beta}}_{i}$. Because the bias remains comparable to KernelSHAP, this geometric variance contraction provides a principled way to decide when to stop iterating.
@@
where $\epsilon$ is a small tolerance (e.g. 0.025). This criterion ensures that the residual uncertainty in any coordinate is a negligible fraction of the overall range of SV. To track $\sigma_{i}^{(n)}$ online with negligible extra memory, we employ Welford’s algorithm Welford01081962 within Algorithm 1.
@@
Since $10<t<$ 1 and $10<\tfrac{\lambda}{\alpha+\lambda}<$ 1, $\rho=t\,\frac{\lambda}{\alpha+\lambda}$ satisfies $\rho<1$, guaranteeing linear convergence of SIM-Shapley.
Theorem 2 reveals a fundamental trade-off between convergence speed and bias, governed by the choice of $t$ and $\lambda$. The detailed proof is in Appendix A.2. In practice, our objective is accurate SV approximation rather than exact fixed-point attainment, so we continue to use the variance-based stopping rule of Eq. (11) for early termination.
4 Stability Improvement
In this section, we introduce two techniques to enhance the convergence stability of SIM-Shapley. By combining these methods, we propose Stable-SIM-Shapley (see Appendix C for complete pseudocode), an improved variant with more robust convergence behavior.
4.1 Preventing Negative Sampling
and if
$r\;>\;\xi,\quad 0<\xi<1,$
we reject the batch and resample. In our experiments, $\xi\in\{0.30,0.35\}$ works well.
Discarding a single batch incurs only the marginal cost of one additional mini-batch draw, since we already track variances for convergence detection. This simple check dramatically reduces the frequency of high-variance updates, yielding smoother, faster convergence in practice.
4.2 Correcting Initialization Bias
To address this, we initialize $\bm{\beta}^{(0)}=0$ and apply a bias correction technique similar to that used in the Adam optimizer kingma2017adammethodstochasticoptimization . This leads to a dynamically adjusted update rule, where the weight factor $t$ varies over iterations. The closed-form solution in Eq. (9) can be adjusted accordingly due to linearity (see Appendix C). The bias-corrected update rule is given by:
@@ 4.2 Correcting Initialization Bias
5:     Solve Eq. (7) to obtain $\bm{\delta}^{(n)}$ and update $\bm{\beta}^{(n)}$

### v2

@@ Abstract
Shapley values provide a principled framework for feature attribution in machine learning, but their exponential computational complexity limits practical applications. We propose SIM-Shapley (Stochastic Iterative Momentum for Shapley values), a novel approximation method that reformulates Shapley value computation as a stochastic optimization problem. Our key insight is that momentum-based updates with adaptive mini-batch sampling can significantly accelerate convergence while maintaining theoretical guarantees. We prove that SIM-Shapley achieves linear $Q$-convergence—a stronger guarantee than existing methods—and derive tight variance bounds under mild assumptions. Three stability mechanisms ensure robust performance: $\ell_{2}$ regularization, negative-sampling detection, and initialization bias correction. Empirically, SIM-Shapley reduces computation time by up to 85% while maintaining or improving attribution quality. Unlike methods requiring specific game formulations, our framework is game-agnostic and extends to general sample average approximation problems. Code is publicly available at https://github.com/nliulab/SIM-Shapley.
1 Background
The deployment of machine learning in high-stakes domains (Ning et al., 2024) demands not only accurate predictions but also interpretable explanations (Doshi-Velez & Kim, 2017). Among existing interpretability methods for explainable AI (XAI), Shapley Value (SV) (Shapley, 1952) has become a foundational solution (Lundberg & Lee, 2017; Rodemann et al., 2024). SV‑based explanations support contrastive explanations, identify corrupted or influential features, and often align well with human intuition (Covert et al., 2020; Molnar, 2025).
However, computing exact Shapley values requires evaluating $\mathcal{O}(d\cdot 2^{d})$ feature coalitions, rendering them computationally prohibitive even for moderate dimensionality. Despite recent progress (Jethani et al., 2022; Zhang et al., 2024; Musco & Witter, 2025) in accelerating SV computation, these works remain fundamentally limited: they lack transparency (introducing additional black-box interpreters), generalizability (requiring specific game/imputer configurations), and completeness in implementations (supporting only local or global explanations).
To fill the gap, we introduce SIM-Shapley, a stochastic optimization framework that fundamentally re-conceptualizes SV computation. Rather than constraining the method to specific games or architectures, we reformulate SV estimation by iteratively refining estimates via mini-batch sampling and momentum updates (Ruppert, 1988).
A high-level comparison of SIM-Shapley with contemporary SV methods is presented in Table 1, demonstrating the advantages of our proposed method. As shown, SIM-Shapley uniquely provides both game and imputer agnosticism—capabilities absent in existing methods—while maintaining minimal hyperparameter overhead and full interpreter transparency.
Our main contributions include: (1) A new SV estimator with provable linear $Q$-convergence, providing faster convergence than state-of-the-art methods while retaining consistency; (2) Stability mechanisms: $\ell_{2}$ regularization, detection of negative-sampling events, and correction of initialization bias and (3) A scalable framework compatible with diverse cooperative game formulations, naturally integrating techniques such as paired sampling (Covert & Lee, 2021), and extensible to a wider class of sample‑average‑approximation (SAA) objectives.
Table 1: Comparison of SIM-Shapley with state-of-the-art SV-based XAI methods. ✓ indicates full support, ✗ indicates no support.
Property FastSHAP SimSHAP LeverageSHAP SIM-Shapley
(Jethani et al., 2022) (Zhang et al., 2024) (Musco & Witter, 2025)
Local/Global Explanations ✓/✗ ✓/✗ ✓/✗ ✓/✓
Classification/Regression Task ✓/✗ ✓/✗ ✗/✓ ✓/✓
Interpreter Transparency
Hyperparameter Overhead High High Minimal Minimal
Game Agnostic
Imputer Agnostic
Public Code Available
1.1 Preliminaries
1.1.1 Shapley Value as Weighted Least Squares
@@
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ denote the $d$ features for $n$ samples, and let $D=\{1,2,\dots,d\}$ represent their index set. Given a predictive model $f$ for the response $Y$, the performance of any subset $S\subseteq D$ can be expressed through a cooperative game $v(S)$. A restricted model $f_{S}=f(x_{S},X_{S^{c}})$ uses only features in $S$, with the complement $S^{c}=D\setminus S$ drawn from the conditional distribution $p(X_{S^{c}}\mid X_{S}=x_{S})$, yielding local explanations.
@@
Several choices for $v(S)$ exist. SHAP (Lundberg & Lee, 2017) defines $v(S)$ as the expectation of the prediction from $f_{S}$, whereas Shapley Effects (Owen, 2014) use output variance. In this work, we adopt the prediction-loss game:
@@
following SAGE (Covert et al., 2020) unless otherwise specified. Besides local explanations, global explanations arise by treating an input-label pair $U=(X,Y)$ as exogenous and defining a stochastic cooperative game as:
@@
which captures the expected contribution of a feature subset across the entire data distribution (Covert et al., 2020). This formulation provides a unified framework that encompasses both local and global SV-based methods. The Shapley value (SV) $\phi_{i}$ of feature $i\in D$ is defined as the average marginal contribution of feature $i$ across all subsets $S$ that exclude $i$:
@@
Lundberg & Lee (2017) recast Shapley estimation as kernel‑weighted linear regression, later refined by Covert & Lee (2021). Approximating $v(S)$ with the additive form $u(S)=\beta_{0}+\sum_{i\in S}\beta_{i}$ leads to the weighted least squares objective:
@@
where $\mu_{\text{Sh}}(S)$ is the Shapley kernel that yields the optimal solution equal to SV (Charnes et al., 1988; Lundberg & Lee, 2017). Specifically, both the empty set $\emptyset$ and the superset $D$ lead to $\mu_{\text{Sh}}(S)$ diverging to infinity. To satisfy the minimization objective in equation 4, the squared term must be forced to vanish, which introduces the constraint condition stated in equation 5.
Let $\bm{\beta}=(\beta_{1},\dots,\beta_{d})$ denote the optimal coefficients (i.e., the minimizer) for equation 4; these are the estimated SV. To represent subsets more conveniently, map each subset $S\subseteq D$ to a binary vector $z\in\{0,1\}^{d}$, with $z_{i}=1\ \Leftrightarrow i\in S$. With a slight abuse of notation, write $v(z)=v(S)$ and $\mu_{\text{Sh}}(z)=\mu_{\text{Sh}}(S)$ for the subset $S=\{i:z_{i}=1\}$. We then define a sampling distribution $p(z)\propto\mu_{\text{Sh}}(z)$ over binary vectors $z$ satisfying $0<\mathbf{1}^{\top}z<d$; otherwise, we set $p(z)=0$.
Under this formulation, estimating SV reduces to solving the following constrained optimization problem (Covert & Lee, 2021):
@@
1.1.2 Monte Carlo Estimation and Practical Challenges
Evaluating equation 5 on all $2^{d}$ subsets is computationally prohibitive. Instead, we draw $m$ i.i.d. samples $z_{i}\sim p(z)$ and their values $v(z_{i})$, yielding the Monte‑Carlo surrogate:
@@ 1.1.2 Monte Carlo Estimation and Practical Challenges
Covert & Lee (2021) derive a closed-form solution via the Karush–Kuhn–Tucker (Boyd & Vandenberghe, 2004) (KKT) conditions and proved consistency as $m\to\infty$. In practice, however, only a finite number of $z_{i}$ are available, which may lead to a singular system in certain regions of the solution space and cause the optimization to collapse. Moreover, naively averaging batch estimates yields high variance and fluctuation due to insufficient coverage of the subset space, leading to numerical instability and optimization failure.
1.1.3 From Sample Average Approximation to Stochastic Iteration
The constrained problem in equation 5 and its sampled version in equation 6 follow the Sample Average Approximation (SAA) paradigm (Shapiro & Wardi, 1996): we replace the expectation in equation 5 by the empirical average in equation 6. Under standard convexity and regularity conditions, the SAA solution converges almost surely as $m\to\infty$ (Fu, 2006; Kim et al., 2015). However, large $m$ can be impractical due to computational constraints, and not all samples are equally informative for improving the estimator.
Therefore we turn to Stochastic Approximation (SA) methods (Hannah, 2015; Shapiro, 1996), which update the solution incrementally using only a mini-batch at each step. This avoids repeatedly solving the full SAA problem and can lead to faster convergence in practice. As pointed out by Royset & Szechtman (2013), SA can outperform SAA in convergence rate, regardless of whether the underlying solver is sublinear, linear, or superlinear.
@@ 2.1 Stochastic Iteration with Momentum
Following the stochastic optimization formulation discussed in Section 1.1.3, we reformulate the KernelSHAP approximation (Lundberg & Lee, 2017) in equation 6 as an iterative procedure to update estimates $\bm{\beta}^{(n)}$ using an exponential moving average (EMA) scheme:
@@
where $0<t<1$ is a fixed momentum parameter. $\bm{\delta}^{(n)}$ represents the sampling information obtained at the $n$-th iteration, which can be interpreted as a form of momentum. The updated estimator $\bm{\beta}^{(n)}$ is constructed as a weighted combination of the previous estimate $\bm{\beta}^{(n-1)}$ and $\bm{\delta}^{(n)}$. In addition, $\bm{\beta}^{(n)}$ shall satisfy the constraint in equation 5 to preserve the additive property of SV, as reflected in the definition of $u(S)$ (equation 4). Although $\bm{\delta}^{(n)}$ can also be adopted as the optimization variable owing to its linearity with $\bm{\beta}^{(n)}$, the optimization in equation 7 is performed with respect to $\bm{\beta}^{(n)}$ to preserve notational consistency in Section 1.1.
@@
For notational simplicity, we do not distinguish between $v(S)$ in equation 1 and $\mathbf{V}(S)$ in equation 2, as replacing $v(S)$ with $\mathbf{V}(S)$ allows $\bm{\beta}^{(n+1)}$ in SIM-Shapley to be interpreted as an estimator of the global SV explanation. To approximate $\mathbf{V}(S)$, we construct a Global Reference Set by sampling from the original dataset, thereby approximating the distribution $p(U)$ and enabling the evaluation of $\mathbb{E}_{U}[v(S,U)]$ for each cooperative game.
As discussed in Section 1.1.1, to approximate the conditional distribution $p(X_{S^{c}}\mid X_{S}=x_{S})$, we handle the missing features $X_{S^{c}}$ by marginalizing them using their joint marginal distribution. Specifically, we build a Background Set from the data and use it to impute the removed variables. Since our focus is on estimating the SV rather than the underlying games themselves, this treatment does not affect the convergence or correctness of the iterative procedure.
Building on these ideas, we propose the SIM-Shapley algorithm, together with its local explanation version described in Algorithm 1. The global version is provided in the Appendix C. We apply the KKT conditions to equation 7, leveraging the convexity of the objective function and the linearity of the constraint. Given a sampled batch $(z_{1},\dots,z_{m})$, the associated Lagrangian function, with Lagrange multiplier $\nu\in\mathbb{R}$, is defined as follows:
@@ 2.2 Variance and Convergence
We further analyze the variance and convergence behavior of the SIM-Shapley estimator $\bm{\beta}^{(n)}$ with respect to the ideal SV vector $\bm{\beta}$. Throughout this paper, we treat the KernelSHAP estimator in equation 6 as the ground truth, consistent with prior works demonstrating that KernelSHAP and its variants are either unbiased or nearly unbiased (Covert & Lee, 2021; Covert et al., 2020; Chen et al., 2022; Lundberg & Lee, 2017). Since SIM-Shapley is designed to approximate KernelSHAP through a stochastic iterative scheme with regularization, any potential bias is empirically negligible with a tiny $\lambda$, as confirmed by Figure 1 and our experiments in Section 4.
Figure 1: Bias-regularization trade-off for SIM-Shapley on the Bank dataset. Bias represents the $\ell_{2}$ distance from ground-truth SV as $\lambda$ varies from 0 (no regularization) to 1.
@@
Below we analyze SIM-Shapley’s variance and convergence across iterations, as these determine practical performance in typical settings with negligible bias.
@@
Let $\hat{\bm{\beta}}$ denote the estimator obtained from KernelSHAP in equation 6 with $m$ samples of $z$, and let $\bm{\beta}^{(n)}$ be the SIM-Shapley iterates with fixed momentum $t\in(0,1)$. As $m\to\infty$, for fixed momentum $t\in(0,1)$ and any feature index $i$ and any iteration $n\geq 1$,
@@
A detailed proof of Theorem 1 is given in Appendix B.1. In essence, each EMA update in SIM-Shapley reduces the variance of feature $i$ by at least a factor of $(1-t)^{2}$ relative to the vanilla SAA estimator $\hat{\bm{\beta}}_{i}$. Because the bias remains comparable to KernelSHAP, this geometric variance contraction provides a principled way to decide when to stop iterating.
@@
where $\epsilon$ is a small tolerance (e.g. 0.025). This criterion ensures that the residual uncertainty in any coordinate is a negligible fraction of the overall range of SV. To track $\sigma_{i}^{(n)}$ online with negligible extra memory, we employ Welford’s algorithm (Welford, 1962) within Algorithm 1.
@@
Since $0<t<1$ and $0<\tfrac{\lambda}{\alpha+\lambda}<1$, $\rho=t\,\frac{\lambda}{\alpha+\lambda}$ satisfies $\rho<1$, guaranteeing linear convergence of SIM-Shapley.
Theorem 2 reveals a fundamental trade-off between convergence speed and bias, governed by the choice of $t$ and $\lambda$. The detailed proof is in Appendix B.2. In practice, our objective is accurate SV approximation rather than exact fixed-point attainment, so we continue to use the variance-based stopping rule of equation 11 for early termination.
@@
5:  Solve equation 7 to obtain $\bm{\delta}^{(n)}$ and update $\bm{\beta}^{(n)}$
@@
3 Stability Improvement
In this section, we introduce two techniques to enhance the convergence stability of SIM-Shapley. By combining these methods, we propose Stable-SIM-Shapley (see Appendix D for complete pseudocode), an improved variant with more robust convergence behavior.
3.1 Preventing Negative Sampling
and if $r\;>\;\xi,\quad 0<\xi<1$, we reject the batch and resample. In our experiments, $\xi\in\{0.30,0.35\}$ generally works well. This incurs minimal cost: one additional mini-batch draw, since variances are already tracked for convergence detection. This check significantly reduces high-variance updates, improving convergence.
3.2 Correcting Initialization Bias
To address this, we initialize $\bm{\beta}^{(0)}=0$ and apply a bias correction technique similar to that used in the Adam optimizer (Kingma & Ba, 2017). This leads to a dynamically adjusted update rule, where the weight factor $t$ varies over iterations. The closed-form solution in equation 9 can be adjusted accordingly due to linearity (see Appendix D). The bias-corrected update rule is given by: